OpenStudy (zasharra):

I really need help, and i don't have much time. plz somebody i'm not just asking for answers .. I AM WILLING TO WORK!! help plz

OpenStudy (anonymous):

lol ok whats the question

OpenStudy (zasharra):

ok i did part on but i have no clue how to do the rest.

OpenStudy (anonymous):

ok

OpenStudy (zasharra):

2.Draw another right triangle. Roll a die one time and record the number you roll. Label one of the legs of a right triangle with the number you rolled. Now, roll the die two more times and record the numbers you roll. Use the two numbers you rolled to create an angle measure for your triangle. On the drawing of your right triangle, label the leg and angle of the triangle using the numbers you rolled. Find the length of the missing leg and find the missing angle measure. Write a paragraph telling how you found the missing values. You must show all work to receive credit. Here's an example below. Notice the dice have values above the values you will have on your six-sided dice. Please use your own original values from rolling your dice: Roll # 1: A die is shown with the number 11 visible. Rolls # 2 and # 3: A picture of two dice is shown. One die is showing the number 7 and the other die is showing the number 9. If an eleven is rolled, then label one leg of the triangle eleven units in length. If the next two rolls yield a seven and nine, then label one angle as 79°. A right triangle is shown. One leg is labeled 11 and an interior angle is labeled 79 degrees. 3.Draw another right triangle. Roll a die two times and record the numbers you roll. On the drawing of your right triangle, label the legs of the triangle using the numbers you rolled. Use the Pythagorean Theorem and trigonometric functions to find the length of the hypotenuse and the missing angle measures. Write a paragraph telling how you found the missing values. You must show all work to receive credit. Here's an example below. Notice the dice have values above the values you will have on your six-sided dice. Please use your own original values from rolling your dice: Rolls # 1 and # 2: Two dice are shown. One die is showing the number 9 and the other die is showing the number 11.

OpenStudy (anonymous):

ok what do you have so far?

OpenStudy (zasharra):

The teacher want just a brief description of what i did so here is my answer for the first part: 1. The numbers that were rolled are 6 and 5.  I drew my right triangle, and the Pythagorean Theorem was 6^2 + 5^2 = c^2 (a^2 + b^2 = c^2)  Once I solved that, I got 61  So length of the hypotenuse would be square root of 61 (7.8)

OpenStudy (zasharra):

@billj5

OpenStudy (anonymous):

ok so that looks good now just put what you have there in paragraph form and you're good to go

OpenStudy (zasharra):

ok. dont know what to do on the rest of the question

OpenStudy (anonymous):

oh the angle part, sorry

OpenStudy (zasharra):

1. a) The numbers that were rolled are 6 and 5. b) I drew my right triangle, and the Pythagorean Theorem was 6^2 + 5^2 = c^2 (a^2 + b^2 = c^2) c) Once I solved that, I got 61 d) So length of the hypotenuse would be square root of 61 (7.8)

OpenStudy (zasharra):

thats my answer for part one

OpenStudy (zasharra):

@billj5 r u there?

OpenStudy (anonymous):

label the angles x and y

OpenStudy (anonymous):

now use the trig identities to find the angles, you know one of the angles is 90 degrees

OpenStudy (anonymous):